OpenStudy (anonymous):

Using the product rule to differentiate the following: Y = 8sin(5x)loge(5x)

OpenStudy (anonymous):

\[(fg)'=f'g+g'f\] with \[f(x)=\sin(5x), g(x)=\ln(5x)\] so the first thing you need to find is \(f'(x)\) and \(g'(x)\) . both require the chain rule for the first one since \(\frac{d}{dx}\sin(x)=\cos(x)\) you have \(\frac{d}{dx}\sin(5x)=5\cos(5x)\)

OpenStudy (anonymous):

so let y=u*v and let u=8sin(5x) v=loge(5x) differentiate your 'u' to get u' and v to get v' then the formula for the derivative = v*u' + u*v' i'm a little unclear as to whether it is log(5x) or log(e^5x) if you could clarify then i can confirm the answer you get

OpenStudy (anonymous):

I'm here but I was writing down my own answer

OpenStudy (anonymous):

This was my answer : dy/dx= 8sin5x/x + 8cos5xloge5x

OpenStudy (anonymous):

Did I get it right?

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

oops no, hold the phone!

OpenStudy (anonymous):

this part 8cos5xloge5x is wrong

OpenStudy (anonymous):

Because in the textbook the answer was dy/dx= 8sin5x/x + 40cos5xloge5x

OpenStudy (anonymous):

if \(f(x)=\sin(5x)\) then \(f'(x)=5\cos(5x)\) so you are missing a factor of 5 in that part

OpenStudy (anonymous):

it is easy to forget the chain rule for something like \(\sin(5x)\) but you need a factor of 5 out front

OpenStudy (anonymous):

I almost got it right :) but thanks for the help

OpenStudy (anonymous):

you don't see the factor in the derivative of \(\ln(5x)\) because \(\ln(5x)=\ln(5)+\ln(x)\) and \(\ln(5)\) is just a constant

OpenStudy (anonymous):

Yes Chad that answer is correct

OpenStudy (anonymous):

dy/dx= 8sin5x/x + 40cos5xloge5x is correct :)